On this handout, any word in a program line that begins with an uppercase letter is actually a command and must be retrieved from a menu or key. For example:

X is X,T,θ,n

Y1 is VARS then Y-VARS then 1: Function then 1: Y_{1}

→ is STO⇒

Disp is PRGM then I/O then 3:Disp IS>( is PRGM then CTL then A:IS>